2. A publishing house hed 40% of its total sales in Math books, 10% in pupil diary, 20%

in Filipino books, 20% in Science books and 10% in English books. If the sales

amounted to P1,250,000.00 find the total sales in each kind of book. What book had

the least sales

Solution:


1
Expert's answer
2021-06-21T06:47:06-0400
"40+10+20+20+10=100(\%)"


"\begin{matrix}
 1250000.00 & 100\% \\
 x & 1\%
\end{matrix}"

"x=\dfrac{1250000.00(1)}{100}=12500.00"

Math


"12500.00(40)=500,000.00"

Pupil diary


"12500.00(10)=125,000.00"



Filipino


"12500.00(20)=250,000.00"



 Science 


"12500.00(20)=250,000.00"



English 


"12500.00(10)=125,000.00"



Pupil diary books and English books had the least sales.